OpenStudy (naveenbhatia1312):

Two triangles can be formed with the given information. Use the Law of Sines to solve the triangles. B = 29°, b = 26, c = 28 A = 92.5°, C = 58.5°, a = 12.6; A = 87.5°, C = 121.5°, a = 12.6 A = 119.5°, C = 31.5°, a = 14.5; A = 2.5°, C = 148.5°, a = 14.5 A = 92.5°, C = 58.5°, a = 53.6; A = 87.5°, C = 121.5°, a = 53.6 A = 119.5°, C = 31.5°, a = 46.7; A = 2.5°, C = 148.5°, a = 2.3

OpenStudy (faiqraees):

\[(\sin A)/a = (\sin B)/b=(\sin C)/c\]

OpenStudy (naveenbhatia1312):

im just starting this...can you explain please

OpenStudy (michele_laino):

using the Law of SInes we can write this: \[\Large \frac{b}{{\sin \beta }} = \frac{c}{{\sin \gamma }} \Rightarrow \frac{{26}}{{\sin 29}} = \frac{{28}}{{\sin \gamma }}\] 

OpenStudy (faiqraees):

All triangles follow the Sine Law. Which is to say that the ratio of an angle and the side with which the angle isn't connected with (the opposite side) is equal to ratio of an angle and its opposite side. Hence (sinA)/a=(sinB)/b=(sinC)/c Plug the values of variables into their corresponding positions and you will work out the values of the unkwown variables.

OpenStudy (michele_laino):

therefore we have: \[\Large \sin \gamma = \frac{{28}}{{26}} \cdot \sin 29 = ...?\]

OpenStudy (naveenbhatia1312):

31.2

OpenStudy (michele_laino):

I got a different result: \(\Large \gamma \simeq 31.5\) degrees

OpenStudy (naveenbhatia1312):

oh wait yeah. my mistake

OpenStudy (michele_laino):

now, since the sum of interior angles has to be equal to 180 degrees, we can write: \[\Large \begin{gathered} \alpha = 180 - \left( {\beta + \gamma } \right) = \hfill \\ \hfill \\ = 180 - \left( {29 + 31.5} \right) = ...? \hfill \\ \end{gathered} \]

OpenStudy (naveenbhatia1312):

119.5

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

finally, if we apply the Law of Sines again, we can write this: \[\Large \frac{a}{{\sin 119.5}} = \frac{{26}}{{\sin 29}}\]

OpenStudy (michele_laino):

therefore: \[\Large a = \sin \left( {119.5} \right) \cdot \frac{{26}}{{\sin 29}}=...?\]

OpenStudy (naveenbhatia1312):

107.14

OpenStudy (michele_laino):

hint: we have this: \[\Large \begin{gathered} a = \sin \left( {119.5} \right) \cdot \frac{{26}}{{\sin 29}} = \hfill \\ \hfill \\ = 0.870 \cdot \frac{{26}}{{0.485}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (naveenbhatia1312):

46.63

OpenStudy (michele_laino):

that's right!

OpenStudy (naveenbhatia1312):

would the answer be the last choice?

OpenStudy (michele_laino):

yes! It is: A = 119.5°, C = 31.5°, a = 46.7

OpenStudy (naveenbhatia1312):

b and d have a similar answer how do I tell which one is correct?

OpenStudy (naveenbhatia1312):

oh the a

OpenStudy (naveenbhatia1312):

Thank you so much!

OpenStudy (michele_laino):

it is the option of the fourth row, of the left column